# A Large Deviations Perspective on Policy Gradient Algorithms

### v2

@@ Abstract
$\leq\tfrac{1}{2}L_{1}\|\theta_{t}-\theta^{\star}\|_{2}^{2}+\frac{1}{2\mu(\theta_{t})}\|g(\theta_{t})\|_{2}^{2}$
$\leq\tfrac{1}{2}L_{1}\|\theta_{t}-\theta^{\star}\|_{2}^{2}+\frac{L_{1}^{2}}{2\mu(\theta_{t})}\|\theta_{t}-\theta^{\star}\|_{2}^{2}.$
As such, one can set $\gamma_{t}=(1+\eta_{t}L_{1}+\eta_{t}L_{1}^{2}/M\mu(\theta_{t})+\eta_{t}^{2}L_{1}^{2})^{1/2}\leq(3+L_{1}/\mu)^{1/2}=\bar{\gamma}$ such that
$\|\theta_{t+1}-\theta^{\star}\|_{2}$ $\leq\gamma_{t}\|\theta_{t}-\theta^{\star}\|_{2}+\eta_{t}\|Z_{t}\|_{2}\quad t=1,\dots,T.$ (4.3)
Step 4, $\theta\in B^{c}_{\ell}(\theta^{\star},\delta)$. Using Step 3 and Assumption 2.3 one can show that $\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)\leq\mathrm{exp}\left(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2}\right)\mathrm{exp}(\bar{\gamma}(\ell+t_{0}+1)\|\lambda\|_{2}\|\theta-\theta^{\star}\|_{2}).$ Hence, we have
@@
Exploiting Lemma 3.2i.e. $\nu_{\ell}$ being induced by $\mathbb{P}\left(\|\theta_{\ell}-\theta^{\star}\|_{2}\geq\delta\right)\leq e^{1-(\ell+t_{0})L_{1}\delta^{2}/(2\mu K)}$, we know there is a sufficiently large $\delta(\lambda)=O(\bar{\gamma}\|\lambda\|_{2}\mu K/L_{1})$, denoted as $\bar{\delta}(\lambda)$ such that
$\Gamma_{\ell+1|B^{c}_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})\leq\bar{C}\mathrm{exp}\left(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2}\right)$ (4.4)
for some constant $\bar{C}$ cf. [BJK23, Proof of Lem. 6]. Step 5, limiting recursion. Combining (4.2) and (4.4) then yields
$\Gamma_{\ell+1}(\zeta_{\ell})\leq\bar{C}\mathrm{exp}\left(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2}\right)+\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta^{\star})+r(\lambda)\right)\Gamma_{\ell}(\zeta_{\ell-1}),$
where $r(\lambda)=r_{0}(\lambda,\bar{\delta}(\lambda))$. Continuing with the recursion provides us with
@@
Note that $\Gamma_{1}(\zeta_{1})$ is finite by Assumption 2.1. Then we immediately obtain that
@@
To complete the proof, we appeal to [BJK23, Lemma C.3] and find that
@@
Suppose that Assumption 2.12.2, 2.3 and 4.1 hold and set $\eta_{t}=\eta/(t+t_{0}+1)$ as in Lemma 3.2. Define $\Psi$ as in Theorem 4.2. Then, with probability at least $61-\epsilon/$ 6 we have that the sequence $(\theta_{t})_{t}$ satisfies a LDP upper bound with a rate function $I$. The function $I$ is the Legendre-Fenchel transformation of $\Psi$. That is, for any Borel set $\Theta\subseteq\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, we have that
$\limsup_{t\to+\infty}\frac{1}{t}\log\mathbb{P}(\theta_{t}\in\Theta)\leq-\inf_{\theta^{\prime}+\theta^{\star}\in\overline{\Theta}}I(\theta^{\prime}).$ (4.5)
Proof.
Directly from Theorem 4.2 and the Gärtner-Ellis theorem [Gär77, Ell84]. ∎
Note, one can also bring the offset term $\theta^{\star}$ into the left-hand-side of (4.5), i.e., consider $\mathbb{P}(\theta_{t}-\theta^{\star}\in\Theta)$. Indeed, strictly speaking, the LDP upper bound is derived for the sequence $(\theta_{t}-\theta^{\star})_{t}$.
@@ 5   Ramifications
One could argue that the results (bounds) from above contribute towards a better understanding of the softmax transformation, the effect of the regularization parameter $\tau$, the effect of initialization and the analysis of restrictive policy parametrizations, all active topics of research [Hen+20, Mei+20, Li+21, Mei+20a]. However, more interesting is that Theorem 4.3 also allows for a simple proof of exponential decay, with high probability, for a whole family of policy parametrizations different from softmax.
The structure of $\Psi$ (non-negativity and finiteness) immediately reveals that its Fenchel conjugate (Legendre-Fenchel transform) $I$ satisfies $I(0)=0$, better yet, $I(\theta^{\prime})>0$ for all $\theta^{\prime}\neq 0$. Hence, $I$ can be identified as a good rate function [DZ09, p. 4]. Moreover, for any Borel set $\Theta$ such that $\theta^{\star}\not\in\overline{\Theta}$ we have, with high probability, that the probability $\mathbb{P}(\theta_{t}\in\Theta)$ decays exponentially fast. Not only that, we can provide a bound on the convergence rate for any of those sets, i.e., let $r:=\inf_{\theta^{\prime}+\theta^{\star}\in\overline{\Theta}}I(\theta^{\prime})$, then, $\mathbb{P}(\theta_{t}\in\Theta)\leq\mathrm{exp}(-rt+o(t))$. To lift this observation to any continuous transformation of the softmax parametrization we need the following principle from the theory of large deviations. To avoid pathological examples, we assume all our sets to be topological Hausdorff spaces.
Theorem 5.1 (Contraction principle [DZ09, § 4.2.1]).
@@
Although we studied a softmax policy, the contraction principle, originally used in [DV76], allows us to extend our LDP results to any continuous transformation $f$ of $\theta_{t}-\theta^{\star}$e.g., one recovers the escort transformation [Mei+20] by transforming $\theta_{t}$ component-wise by the map $\mathbb{R}\ni w\mapsto p\log|w|$ for $p\geq 1$. To be explicit, whereas the softmax policy is given by (2.2), the escort parametrization is given by $\pi_{\theta}(a|s)={|{\theta(s,a)}|^{p}}/{\sum_{a^{\prime}\in\mathcal{A}}|{\theta\left(s,a^{\prime}\right)|^{p}}}$. Hence, we directly infer that the exponential decay rate extends from the softmax- to the escort parametrization. Indeed, $w\mapsto p\log|w|$ is not a smooth function, yet, for $p\geq 2$ the escort policy parametrization is differentiable, which is precisely the setting for which an exponential decay rate under entropic regularization is provided by [Mei+20, Thm. 4]. To formally summarize this observation, by combining Theorem 4.3 and Theorem 5.1 we get the following.
@@
$\limsup_{t\to+\infty}\frac{1}{t}\log\mathbb{P}(\omega_{t}\in\Omega)\leq-\inf_{\omega^{\prime}+\omega^{\star}\in\overline{\Omega}}I^{\prime}(\omega^{\prime}),$ (5.1)
@@
Similar to the escort transformation, one can study spherical- and Taylor softmax [BV16]. However, Corollary 5.2 is ought to be most interesting to apply to prospective policy parametrizations leading to (5.1) being non-trivial. Indeed, we emphasize that one must verify that (5.1) is non-trivial, there is no free lunch.
To remove the $(1-\epsilon/6)$-high probability statements throughout, future work aims at studying the multiphase algorithm from [DZL21] in the context of large deviations, this requires being able to study a controlled sequence $(Z_{t})_{t}$, which is a non-trivial extension. We also remark that in the context of SGD for an objective satisfying a PL condition, the LDP upper bound always holds, simply since $\mu>0$, cf. Lemma 3.1.

### v3

@@ Abstract
$\leq\tfrac{1}{2}L_{1}\|\theta_{t}-\theta^{\star}\|_{2}^{2}+\frac{1}{2\mu(\theta_{t})}\|g(\theta_{t})\|_{2}^{2}\leq\tfrac{1}{2}L_{1}\|\theta_{t}-\theta^{\star}\|_{2}^{2}+\frac{L_{1}^{2}}{2\mu(\theta_{t})}\|\theta_{t}-\theta^{\star}\|_{2}^{2}.$
As such, one can set $\gamma_{t}:=(1+\eta_{t}L_{1}+\eta_{t}L_{1}^{2}/M\mu(\theta_{t})+\eta_{t}^{2}L_{1}^{2})^{1/2}\leq(3+L_{1}/\mu)^{1/2}=:\bar{\gamma}$ such that
$\|\theta_{t+1}-\theta^{\star}\|_{2}$ $\leq\gamma_{t}\|\theta_{t}-\theta^{\star}\|_{2}+\eta_{t}\|Z_{t}\|_{2}\quad t=1,\dots,T.$ (17)
Step 4, $\theta\in B^{c}_{\ell}(\theta^{\star},\delta)$. Using Step 3 and Assumption 3 one can show that
$\allowbreak\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)\leq\mathrm{exp}(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2})\mathrm{exp}(\bar{\gamma}(\ell+t_{0}+1)\|\lambda\|_{2}\|\theta-\theta^{\star}\|_{2}).$ Hence, we have
@@
Exploiting Lemma 3.3, i.e. $\nu_{\ell}$ being induced by $\mathbb{P}\left(\|\theta_{\ell}-\theta^{\star}\|_{2}\geq\delta\right)\leq e^{1-(\ell+t_{0})L_{1}\delta^{2}/(2K)}$, we know there is a sufficiently large $\delta(\lambda)=O(\bar{\gamma}\|\lambda\|_{2}K/L_{1})$, denoted as $\bar{\delta}(\lambda)$ such that
$\Gamma_{\ell+1|B^{c}_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})\leq\bar{C}\mathrm{exp}\left(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2}\right)$ (18)
for some constant $\bar{C}$ cf. (Bajovic et al., 2023, Proof of Lem. 6). Step 5, limiting recursion. Combining (16) and (18) then yields $\Gamma_{\ell+1}(\zeta_{\ell})\leq\bar{C}\mathrm{exp}\left(\tfrac{1}{2}\sigma^{2}\eta^{2}\|\lambda\|_{2}^{2}\right)+\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta^{\star})+r(\lambda)\right)\Gamma_{\ell}(\zeta_{\ell-1}),$ where $r(\lambda)=r_{0}(\lambda,\bar{\delta}(\lambda))$. Continuing with the recursion provides us with
@@
Note that $\Gamma_{1}(\zeta_{1})$ is finite by Assumption 1. Then we immediately obtain that
@@
To complete the proof, we appeal to (Bajovic et al., 2023, Lemma C.3) and find that
@@
The Theorem below is a strict generalization of Lemma 3.3. Specifically, Theorem 4.3 characterizes the concentration rate applicable to any Borel set.
@@
Suppose that Assumption 12, 3 and 4 hold and set $\eta_{t}=\eta/(t+t_{0}+1)$ as in Lemma 3.3. Define $\Psi$ as in Theorem 4.1. Then, with probability at least $61-\epsilon/$ 6 we have that the sequence $(\theta_{t})_{t}$ satisfies a LDP upper bound with a rate function $I$. The function $I$ is the Legendre-Fenchel transformation of $\Psi$. That is, for any Borel set $\Theta\subseteq\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, we have that
$\limsup_{t\to+\infty}\frac{1}{t}\log\mathbb{P}(\theta_{t}\in\Theta)\leq-\inf_{\theta^{\prime}+\theta^{\star}\in\overline{\Theta}}I(\theta^{\prime}).$ (19)
Proof 4.4.
Directly from Theorem 4.1 and the Gärtner-Ellis theorem (Gärtner, 1977; Ellis, 1984).
Note, one can also bring the offset term $\theta^{\star}$ into the left-hand-side of (19), i.e., consider $\mathbb{P}(\theta_{t}-\theta^{\star}\in\Theta)$. Indeed, strictly speaking, the LDP upper bound is derived for the sequence $(\theta_{t}-\theta^{\star})_{t}$.
@@ 5 Ramifications
One could argue that the results (bounds) from above contribute towards a better understanding of the softmax transformation, the effect of the regularization parameter $\tau$, the effect of initialization and the analysis of restrictive policy parametrizations, all active topics of research Hennes et al. (2020); Mei et al. (2020a); Li et al. (2021); Mei et al. (2020b). However, more interesting is that Theorem 4.3 also allows for a simple proof of exponential decay, with high probability, for a whole family of policy parametrizations different from softmax.
The structure of $\Psi$ (non-negativity and finiteness) immediately reveals that its Fenchel conjugate (Legendre-Fenchel transform) $I$ satisfies $I(0)=0$, better yet, $I(\theta^{\prime})>0$ for all $\theta^{\prime}\neq 0$. Hence, $I$ can be identified as a good rate function (Dembo and Zeitouni, 2009, p. 4). Moreover, for any Borel set $\Theta$ such that $\theta^{\star}\not\in\overline{\Theta}$ we have, with high probability, that the probability $\mathbb{P}(\theta_{t}\in\Theta)$ decays exponentially fast. Not only that, we can provide a bound on the convergence rate for any of those sets, i.e., let $r:=\inf_{\theta^{\prime}+\theta^{\star}\in\overline{\Theta}}I(\theta^{\prime})$, then, $\mathbb{P}(\theta_{t}\in\Theta)\leq\mathrm{exp}(-rt+o(t))$. To lift this observation to any continuous transformation of the softmax parametrization we need the following principle from the theory of large deviations. To avoid pathological examples, we assume all our sets to be topological Hausdorff spaces.
Theorem 5.1 (Contraction principle (Dembo and Zeitouni, 2009, § 4.2.1)).
@@
Although we studied a softmax policy, the contraction principle, originally used by Donsker and Varadhan (1976), allows us to extend our LDP results to any continuous transformation $f$ of $\theta_{t}-\theta^{\star}$e.g., one recovers the escort transformation (Mei et al., 2020a) by transforming $\theta_{t}$ component-wise by the map $\mathbb{R}\ni w\mapsto p\log|w|$ for $p\geq 1$. To be explicit, whereas the softmax policy is given by (3), the escort parametrization is given by $\pi_{\theta}(a|s)={|{\theta(s,a)}|^{p}}/{\sum_{a^{\prime}\in\mathcal{A}}|{\theta\left(s,a^{\prime}\right)|^{p}}}$. Hence, we directly infer that the exponential decay rate extends from the softmax- to the escort parametrization. Indeed, $w\mapsto p\log|w|$ is not a smooth function, yet, for $p\geq 2$ the escort policy parametrization is differentiable, which is precisely the setting for which an exponential decay rate under entropic regularization is provided by (Mei et al., 2020a, Thm. 4). To formally summarize this observation, by combining Theorem 4.3 and Theorem 5.1 we get the following.
@@
$\limsup_{t\to+\infty}\frac{1}{t}\log\mathbb{P}(\omega_{t}\in\Omega)\leq-\inf_{\omega^{\prime}+\omega^{\star}\in\overline{\Omega}}I^{\prime}(\omega^{\prime}),$ (20)
@@
Similar to the escort transformation, one can study spherical- and Taylor softmax (de Brébisson and Vincent, 2016). However, Corollary 5.2 is ought to be most interesting to apply to prospective policy parametrizations leading to (20) being non-trivial. Indeed, we emphasize that one must verify that the solution to the non-linear optimization problem (20) is non-trivial, there is no free lunch. Moreover, the gradient estimation errors needs to remain well-behaved after the transformation, which may not be true for an arbitrary continuous mapping $f$. We leave it to future work to explore the exact characterizations of the scope of Corollary 5.2.
To remove the $(1-\epsilon/6)$-high probability statements throughout, future work aims at studying the multiphase algorithm from (Ding et al., 2021) in the context of large deviations, this requires being able to study a controlled sequence $(Z_{t})_{t}$, which is a non-trivial extension. We also remark that in the context of SGD for an objective satisfying a PL condition, the LDP upper bound always holds, simply since $\mu>0$, cf. Lemma 3.1.